OpenStudy (lovesummer731):

Calculate the number of grams of nitrogen gas required to make 1.22 L of ammonia at STP

OpenStudy (ipwnbunnies):

Use the Ideal Gas Equation: PV = nRT; where 'n' is the amount of moles of ammonia, 'R' is the gas constant.

OpenStudy (ipwnbunnies):

At STP, P = 1 atmosphere and T = 293 Kelvin

OpenStudy (ipwnbunnies):

We wants to find how many moles of ammonia are in 1.22 L at STP, for this first step.

OpenStudy (lovesummer731):

so how would you find ammonia? i'm sorry but i am just terrible at chemistry

OpenStudy (ipwnbunnies):

We plug in the numbers we got so far.

OpenStudy (ipwnbunnies):

\[(1 atm)(1.22 L) = n*(0.08206 \frac{L-atm}{mol-K})(293 K)\]

OpenStudy (ipwnbunnies):

R is a gas constant, so it never changes. Solve for 'n', that'll be the amount of moles of ammonia.

OpenStudy (lovesummer731):

ok

OpenStudy (lovesummer731):

can you cross out K because when you solve for "n", you can divide potassium by itself and they cancel out. would that be correct?

OpenStudy (ipwnbunnies):

Err, that K means Kelvin. And yes, all the units should be crossed out, except 'moles'.

OpenStudy (lovesummer731):

oh wait... not potassium, it would be the Kelvins

OpenStudy (lovesummer731):

(1atm)(I.22L)= n(24.04358L-atm/mol) is what i have so far.

OpenStudy (ipwnbunnies):

Ok, now solve for n. It took you like 10 minutes to multiply?

OpenStudy (lovesummer731):

no the atm and the mol was confusing me so i just started over and checked what i had already to make sure i was doing it right

OpenStudy (ipwnbunnies):

The Liters and atmospheres units cancelled out when you divide.

OpenStudy (lovesummer731):

so n= .o507411mols?

OpenStudy (ipwnbunnies):

Not sure, lol. didn't do the math, hold on.

OpenStudy (lovesummer731):

lol ok

OpenStudy (ipwnbunnies):

Yes, so we have 0.0507 moles of ammonia, NH3. Can you use stoichiometry to figure out how many moles of Nitrogen, N, are in the ammonia?

OpenStudy (lovesummer731):

i think so

OpenStudy (ipwnbunnies):

Ok, we have 0.0507 moles of NH3. How many moles of N are there in it?

OpenStudy (lovesummer731):

wouldn't there be 28g of moles in nitrogen?

OpenStudy (ipwnbunnies):

Err, what? The molar mass of Nitrogen is 14 grams per mol.

OpenStudy (ipwnbunnies):

We're not there yet. Did you find how many moles of nitrogen there are?

OpenStudy (lovesummer731):

how did you get 14?

OpenStudy (ipwnbunnies):

It's from the periodic table lol.

OpenStudy (lovesummer731):

oh hehe lol

OpenStudy (ipwnbunnies):

An atom of nitrogen weights 14 Daltons, a mole of nitrogen measures 14 grams.

OpenStudy (lovesummer731):

in NH3, there is only one atom of nitrogen so would that make the mole of nitrogen in NH3 only 14?

OpenStudy (ipwnbunnies):

Ehh, no, take it back. We got 0.0507 moles of NH3. For every mole of NH3, there is 1 mole of N, nitrogen. We get that from the molecular formula. So... \[(0.0507 mol NH_3)(\frac{1 mol N}{1 mol NH_3}) = 0.0507 mol N\]

OpenStudy (ipwnbunnies):

Now, we have 0.0507 moles of N. We can now use the molar mass of N to figure out the mass.

OpenStudy (lovesummer731):

alright so we just solve for N?

OpenStudy (ipwnbunnies):

No, sorry lol. N means nitrogen.

OpenStudy (ipwnbunnies):

I was trying to label both NH3 and N.

OpenStudy (ipwnbunnies):

We multiply the 0.0507 moles of N by the molar mass to get the grams of nitrogen.

OpenStudy (lovesummer731):

0.7098 grams of nitrogen

OpenStudy (ipwnbunnies):

Correct.

OpenStudy (lovesummer731):

awesome! thx!! so is that the answer?

OpenStudy (lovesummer731):

i gtg but thx for helping me so much.